### v3

# Orbital Description of Landau Levels

Abstract
The pursuit of a lattice analogue for Landau levels has been a central theme in condensed matter physics. Although the correspondence between Chern bands and the lowest Landau level has been widely studied, a lattice realization of the first Landau level remains elusive. Here we construct a minimal lattice model that provides a concrete orbital description of both the lowest and first Landau levels. Using maximally localized Wannier functions with $s$, $p_{-}$, and $p_{+}$ orbital character, we develop a three-orbital model in which the two lowest Chern bands are flat and each carries a Chern number $\mathcal{C}=1$. The band topology arises from a sequence of ideal band inversions between Wannier states at the $\Gamma$ and $K$ points in momentum space, establishing an adiabatic connection between the atomic insulator limit and Landau level physics. Notably, many-body exact diagonalization reveals that the non-Abelian state can appear in the half-filled first Chern band. This construction can be further generalized to realize flat Chern bands analogous to higher Landau levels. Our results provide a new perspective on lattice analogues of Landau levels and may enable the exploration of fascinating topological phenomena at elevated temperatures.
Introduction—The recent discovery of Abelian fractional Chern insulators (FCI) at zero magnetic field, dubbed the fractional quantum anomalous Hall state in moiré MoTe2 and pentalayer graphene [1, 2, 3, 4, 5] has generated intense interest in this new state of matter [6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19]. The emergence of fractional topological states is attributed to the existence of flat Chern bands [20, 21, 22, 23, 24, 25, 26] with nearly ideal quantum geometry in moiré superlattice [27, 28, 29, 30, 31], which resembles the lowest Landau level (LLL). The exotic non-Abelian topological orders [32], such as Moore-Read states [33, 34, 35, 36] or Read-Rezayi state [37] are predicted to exist in the partially filled first Landau level (1LL). These states support non-Abelian quasiparticle excitations [34] and could be utilized as a platform for fault-tolerant quantum computation [38, 39]. This raises an interesting question [40, 41, 42, 43, 44, 45]: Is it possible to construct a lattice analogy of 1LL without explicitly invoking the magnetic field? This question is of significant importance because such a construction could pave the way for realizing non-Abelian fractionalization at elevated temperatures.
Here, rather than relying on the operator algebra to bridge Chern bands and LLs, we establish this connection via an adiabatic pathway from atomic limits to LL physics within the skyrmion Chern band model. We construct a projected three-band lattice model using maximally localized Wannier functions (MLWFs) with $s$, $p_{-}$, and $p_{+}$ orbital characteristics. The LL-like band topology in our model emerges through a sequence of ideal band inversions at high-symmetry points. We show using many-body exact diagonalization that the half-filled first Chern band (1CB) supports non-Abelian Moore-Read state. Our approach can be further extended to engineer flat Chern bands that mimic higher LLs. Notably, the corresponding lattice model is well-suited for implementation in cold atom systems, where the long-range Coulomb interaction is effectively projected onto a short-range density–density interaction.
Intuition—As a starting point for constructing Chern bands in a lattice model, it is essential to understand the characteristics of the corresponding Wannier states. In earlier work, Qi constructed the one-dimensional MLWF in Chern bands that have a one-to-one mapping to the LLL wave functions in quantum Hall (QH) [23] of the form $\psi_{0,K_{y}}(x,y)\sim e^{iK_{y}y}H_{0}(x^{\prime})e^{-x^{\prime 2}/2l^{2}_{B}}$, where $x^{\prime}\equiv x-K_{y}\ell_{B}^{2}$. $H_{0}(x^{\prime})$ is the Hermite polynomial for $n=0$ LLL, and $H_{0}(x^{\prime})e^{-x^{\prime 2}/2l^{2}_{B}}=e^{-x^{\prime 2}/2l^{2}_{B}}$ is related to MLWF of Chern band and exhibits $s$-orbital–like character in one dimension. Intuitively, when this construction is extended to the 1LL, the MLWFs in the corresponding Chern band map to $H_{1}(x^{\prime})e^{-x^{\prime 2}/2l^{2}_{B}}=2x^{\prime}e^{-x^{\prime 2}/2l^{2}_{B}}$, which has a $p$-orbital–like structure. This suggests that, at a minimum, Wannier states with $s$- and $p$-orbital characteristics are necessary to construct a Chern band that captures the physics of the 1LL.
Model and adiabatic connection—We begin with the skyrmion Chern band model in which itinerant electrons couple to a layer pseudospin skyrmion lattice. This model could well describe the consecutive topologically nontrivial flat bands in twisted homobilayer MoTe2 around twist angle $2^{\circ}$ [41, 46, 45, 47, 48, 49, 50, 51],
$\mathcal{H}_{0}=\frac{\mathbf{p}^{2}}{2m}+\mathcal{J}\bm{\sigma}\cdot\mathbf{S}(\mathbf{r}).$ (1)
Here, $\bm{\sigma}$ denotes the pseudospin operator associated with the layer degree of freedom, and $\mathbf{S}(\mathbf{r})=\mathbf{S}(\mathbf{r}+\mathbf{a}_{1,2})$ represents the periodic moiré potential and is naturally coupled to the pseudospin of the layer with strength $\mathcal{J}$. In the small twist-angle limit, the moiré potential becomes dominant, such that $\mathcal{J}\gg\hbar^{2}/ma^{2}$, and enforces the local alignment of the electron layer polarization with $\mathbf{S}(\mathbf{r})$. This alignment induces a pseudospin Berry phase. The effect can be made explicit through a position-dependent $SU(2)$ unitary transformation $\mathbf{U}(\mathbf{r})$, which rotates the pseudospin texture $\mathbf{S}(\mathbf{r})$ into $S(\mathbf{r})\hat{z}$, thereby introducing a gauge field $\mathbf{A}_{j}=(i\hbar/e)\mathbf{U}^{\dagger}\partial_{j}\mathbf{U}$. In the large-$\mathcal{J}$ limit, the system can be projected onto the low-energy manifold of locally layer-polarized electrons, yielding an effective Hamiltonian for the low-energy band as
$\mathcal{H}_{\text{eff}}=\frac{\left(\mathbf{p}-e\mathbf{A}(\mathbf{r})\right)^{2}}{2m}+\sum\limits_{j=x,y}\frac{\hbar^{2}}{8m}(\partial_{j}\hat{\mathbf{S}})^{2}-\mathcal{J}S(\mathbf{r}),$ (2)
[…]
From such observations, we see that $\mathcal{H}_{\text{eff}}$ actually links effective atomic orbital and LL physics in the low-energy manifold. This is made clearly by calculating the band structure of $\mathcal{H}_{0}$ when we take $\mathbf{S}(\mathbf{r})=\mathbf{N}(\mathbf{r})/N^{\lambda}(\mathbf{r})$ with
$\mathbf{N}(\mathbf{r})=\frac{1}{\sqrt{2}}\sum_{j=1}^{6}e^{i\mathbf{q}_{j}\cdot\mathbf{r}}\hat{\mathbf{e}}_{j}+N_{0}\hat{z}$, (3)
where $\hat{\mathbf{e}}_{j}=\left(i\alpha\sin\theta_{j},-i\alpha\cos\theta_{j},-1\right)/\sqrt{2}$, $\mathbf{q}_{j}=(4\pi/\sqrt{3}a)\left(\cos\theta_{j},\sin\theta_{j}\right)$, and the angles satisfy $\theta_{2}=\theta_{1}+2\pi/3$, $\theta_{3}=\theta_{1}+4\pi/3$, and $\theta_{j+3}=\theta_{j}+\pi$. This pseudospin texture can be viewed as a sum of three pseudospin spirals forming a triangular skyrmion lattice [55, 56, 57, 58], and the normalization is controlled by $\lambda$. When $\lambda=1$, $S(\mathbf{r})$ becomes spatially uniform, and the low energy bands of $\mathcal{H}_{0}$ are flat Chern bands with $\mathcal{C}=1$, as shown in Fig. 1(e). The lowest two Chern bands are dispersive LL, where their wave functions overlap with the corresponding flat LL exceed $99\%$ [59]. For simplicity, we label them as LLL and 1LL, respectively. Conversely, when $\lambda\rightarrow 0$, the minimum of $\mathcal{J}S(\mathbf{r})$ forms a moiré triangular lattice (see Fig. 1(g)), and three lowest bands in Fig. 1(a) are topologically trivial flat bands resembling the atomic insulator limit, which emerge from $s$, $_{-}, p+p_{+}$ orbital hopping on a triangular lattice.
Figure 1: Band structures and adiabatic connection from atomic insulator to $L$. (a)-(e) The lowest few bands of $\mathcal{H}_{0}$ with different normalization parameter $\lambda$, where the overlap between Bloch state and Wannier function is labeled as different colors, blue for $s$ orbital, green for $p_{-}$ orbital, red for $p_{+}$ orbital and $\mathcal{C}$ denotes the Chern number of lowest three bands. The inset in (c) shows the Brillouin zone (BZ). (f) Band structure of the minimal three-orbital tight-binding model. (g) Real space distribution of $N(\mathbf{r})$. The parameters are $\mathcal{J}/(\hbar^{2}/ma^{2})=52\pi^{2}$, $\alpha=1$ and $N_{0}=0.28$.
Fig. 1 display the evolution of the band structures from flat topologically trivial bands to flat Chern bands. As $\lambda$ changes from 0 to 1, the lowest three bands invert while their $s$, $_{-}, p+p_{+}$ orbital characteristics remain. To see the band inversion clearly, we graphically depict the overlap of the Bloch states with trial Wannier functions of $s$, $p_{-}$, and $p_{+}$ orbitals, symbolized by blue, green, and red, respectively. The lowest three bands first invert with the upper band at $M$ and their Chern numbers become $\mathcal{C}=[0,0,3]$ as shown in Fig. 1(b), then the second and third lowest bands invert at $K$ and switch Chern numbers to $\mathcal{C}=[0,2,1]$ as in Fig. 1(c), finally the lowest and second lowest bands invert at $\Gamma$ and Chern numbers becomes $\mathcal{C}=[1,1,1]$ as in Fig. 1(d,e). The lowest three Chern bands are isolated in Fig. 1(e). The orbital characteristics of the LLL and 1LL here are consistent with the mapping from the hybrid Wannier functions of Chern band to LL wave functions in QH [23].
Wannier projection—Now we demonstrate that $s$, $_{-}, p+p_{+}$ orbitals constitute a complete subspace to describe LLL and 1LL. By constructing the MLWFs [63, 64, 65, 66] from the continuum model, a three-orbital model is developed to describe the lowest two Chern bands in Fig. 1(e), with the Wannier orbitals forming a triangular lattice. The minimal model includes hopping up to the fifteenth nearest neighbor, where its explicit form is in Supplemental Materials [59]. The band structure in Fig. 1(e) does not possess a local gap below which the total Chern number is zero. A set of frozen states is chosen to preserve the topology of the focused Chern bands and the band disentanglement [67] process is then performed to avoid Wannier obstruction [68]. We choose the lowest two Chern bands as frozen states, and get the minimal tight-binding model where only MLWFs with $s$, $_{-}, p+p_{+}$ orbital characteristics are included. The corresponding band structure is calculated in Fig. 1(f). The MLWFs of $s$, $_{-}, p+p_{+}$ orbitals are shown in Fig. 2. The lowest two bands projected from the LLL and 1LL are well reconstructed in Fig. 1(f), which are labeled LCB (lowest Chern band) and 1CB, respectively.
Figure 2: The MLWFs of the lowest few bands in Fig. 1(e) with $s$, $_{-}, p+p_{+}$ orbital characteristics, and the white hexagon labels Wigner-Seitz cell.
Two band geometry indicators are employed to evaluate the different between Chern bands and LL [69, 27, 70, 71, 28, 72, 29, 31], namely Berry curvature fluctuation $\delta\mathcal{B}$ and average trace condition $\mathbb{T}$ (non-negative) defined as
$(\delta\mathcal{B})^{2}$ $\equiv\frac{\Omega_{\text{BZ}}}{4\pi^{2}}\int_{\text{BZ}}d\mathbf{k}\left(\mathcal{B}(\mathbf{k})-\frac{2\pi\mathcal{C}}{\Omega_{\text{BZ}}}\right)^{2},$ (4)
$\mathbb{T}$ $\equiv\int_{\text{BZ}}d\mathbf{k}\left[\mathrm{Tr}(g(\mathbf{k}))\right],$ (5)
where $\mathcal{B}(\mathbf{k})\equiv-2\mathrm{Im}(\eta^{xy})$ is the Berry curvature, $g(\mathbf{k})\equiv\mathrm{Re}(\eta^{\mu\nu})$ is the Fubini-Study metric, and $\eta^{\mu\nu}(\mathbf{k})\equiv\langle\partial^{\mu}u_{\mathbf{k}}|\left(1-|u_{\mathbf{k}}\rangle\langle u_{\mathbf{k}}|\right)|\partial^{\nu}u_{\mathbf{k}}\rangle$ is the quantum geometric tensor, $\mathcal{C}\equiv(1/2\pi)\int d^{2}\mathbf{k}\mathcal{B}(\mathbf{k})$, $\Omega_{\text{BZ}}$ is area of BZ. We plot the distribution of $\mathcal{B}(\mathbf{k})$ and $\text{Tr}[g(\mathbf{k})]$ of the BZ in Fig. 3. For LLL and 1LL, $\mathcal{B}(\mathbf{k})$ remain positive throughout the whole BZ and the distribution is quite homogeneous with relatively small fluctuation $\delta\mathcal{B}$, and $\mathbb{T}$ is almost ideal ($\mathbb{T}=2n+1$ for $n$th flat LL [73]). However, in the tight-binding model, $\mathcal{B}(\mathbf{k})$ is no longer homogenous but concentrates around the band inversion points, and their sign is not always positive (such as in 1CB) throughout the BZ, leading to relatively large fluctuation $\delta\mathcal{B}$ and deviation of $\mathbb{T}$ from the ideal value in LL.
Figure 3: Distribution of Berry curvature $\mathcal{B}(\mathbf{k})$ and trace of Fubini–Study metric $\text{Tr}[g(\mathbf{k})]$ for the two lowest bands of both continuum model and tight-binding model across the Brillouin zone (BZ). The BZ is outlined by dashed hexagon. The bands derived from the continuum model are designated as LLL and 1LL, reflecting their highly ideal character and close resemblance to flat $L$s. In contrast, the bands from the tight-binding model are labeled LCB and 1CB. The Berry curvature fluctuation $\delta\mathcal{B}$ is noticeably larger in LCB and 1CB compared to their LLL and 1LL counterparts, resulting in a stronger deviation of the trace condition measure $\mathbb{T}$ from the ideal value.
These differences between Chern bands of the tight-binding model and dispersive LL of the continuum model are inevitable, since we project out the higher energy degree of freedom in the continuum model and keep only $s$, $_{-}, p+p_{+}$ Wannier states. Then the band inversion among them naturally lead to Berry curvature concentration. Especially for the 1CB, the band inversion between $p_{-}$ and $s$ contributes the negative $\mathcal{B}(\mathbf{k})$ around $\Gamma$ point and the inversion between $p_{-}$ and $p_{+}$ contribute the positive $\mathcal{B}(\mathbf{k})$ around $K$ and $K^{\prime}$ points. Remarkably, we will see that a non-Abelian state can occur even when the Berry curvature fluctuates strongly.
Exact diagonalization—To explore whether the non-Abelian state can appear in 1CB, we now study many-body physics at fractional filling in these Chern bands via numerical diagonalization. To make the many-body calculation tractable, we restrict our variational Hilbert space to that in which $N_{\text{uc}}$ electrons fill the LCB and $N_{e}-N_{\text{uc}}$ electrons remain in the 1CB where $N_{e}$ is the number of electrons. The electron-electron interaction Hamiltonian is obtained by projecting the realistic Coulomb interaction into the MLWFs and keeping the leading terms [59], which is defined as
$\mathcal{H}_{\text{int}}=\frac{U}{2}\sum_{i,a\neq b}n_{i,a}n_{i,b}+V\sum_{\left<ij\right>,a,b}n_{i,a}n_{j,b},$ (6)
where $a$,$b$ = ($s$, $_{-}, p+p_{+}$), $\left<ij\right$> means nearest-neighbor, $U$ and $V$ are the strength of onsite and nearest-neighbor interaction. We adopt an isotropic approximation, neglecting minor variations in the interactions between different orbital pairs [59]. We also neglect the kinetic energy since these two lowest Chern bands are quite flat.
Figure 4: Exact diagonalization and PES for fractional filled CB. (a) Low energy many-body energy spectrum for $1/3$ filled LCB with $N_{\text{uc}}=24$. (b) PES with $N_{A}=3$ for the three degenerate ground states in (a), where $\rho=(1/d)\sum_{i}^{d}\left|\Psi_{i}\right>\left<\Psi_{i}\right|$. $d$ is the degeneracy of ground states. (c), (d) Energy spectrum for $1/2$ filled 1CB with $N_{\text{uc}}=26$ and $N_{\text{uc}}=28$. (e), (f) PES with $N_{A}=3$ for the six degenerate ground states in $N_{\text{uc}}=20$ and $N_{\text{uc}}=28$. Here we only show the lowest energy per momentum sectors in addition to the degenerate ground state.
Fig. 4(a) displays the many-body spectra at filling $\nu=1/3$ (that is, $1/3$ filling of LCB) as a function of crystal momentum $\mathbf{k}=k_{1}\mathbf{T}_{1}+k_{1}\mathbf{T}_{2}$, which is labeled $k=k_{1}+N_{1}k_{2}$. Here $k_{i}=0,...,N_{i}-1$ ($i=1,2$) for system size $N_{\text{uc}}=N_{1}\times N_{2}$ with filled particle number $N_{e}=\nu N_{\text{uc}}$ and $\mathbf{T}_{i}$ are basis vectors of crystal momentum. The cluster size is chosen as $N_{\text{uc}}=4\times 6$, and $U=1$ and $V=0$. There are 3 nearly degenerate ground states well separated by a sizable energy gap from excited states. The approximate ground state degeneracy matches with the expected topological degeneracy of a fractional QH state on a torus. We further calculated different cluster sizes with all other parameters fixed and found that the gap remains, indicating its existence in the thermodynamic limit [59]. The lattice momenta of the degenerate ground states have linear indices $(0,8,16)$ and are in precise agreement with the generalized Pauli principle, which is the hallmark of FCI at $1/3$ filling [25]. To further confirm and distinguish FCI and other competing phases, we subsequently calculated the particle entanglement spectrum (PES) which encodes the information of the quasihole excitations [74, 75], by dividing the whole system into $N_{A}=3$ and $N_{e}-N_{A}$ particles. As shown in Fig. 4(b), we find that there is a clear entanglement gap separating the low-lying PES levels from higher ones for degenerate many body ground states. The number of PES levels below the gap exactly matches the typical counting of quasiparticle excitations resulting from the generalized Pauli principle of $1/3$ Laughlin state. These numerical results suggest that LCB resembles LLL.
Fig. 4(c,d) show the many-body spectra at filling $\nu=3/2$ (that is, $1/2$ filling of 1CB assuming full spin polarization). We assume the energy separation between LCB and 1CB is much larger than the interaction ($U,V$), which prevents electrons in the filled LCB from gaining enough energy to be excited and participate in the many-body interactions involving the partially filled 1CB. Under these assumptions, the exact diagonalization is performed on finite-size torus with $U=1.3$ and $V=1$. For $N_{\text{uc}}=2\times 13$, the number of electrons occupying the 1CB is odd (13), while for $N_{\text{uc}}=2\times 14$, it is even (14). In these two cases, we observe two-fold and sixfold ground state quasi-degeneracies in Fig. 4(c) and 4(d), respectively. These are precisely the degeneracies expected for a Moore-Read state on the torus due to an even-odd effect [76, 77, 34], and the enhancement of gap indicates its existence in the thermodynamic limit. The lattice momenta at which these ground states occur also match the momenta of non-Abelian $\nu$ = 1/2 FCI based on the fractional QH-FCI folding scheme [25, 78]. We also find that under the twist boundary condition, the quasi-degenerate ground states remain well separated from the other low-energy excitation spectrum, indicating the robustness of the excitation gap [59].
To exclude other competing phases in the half-filled 1CB, we further calculate PES with $N_{A}=3$ for $N_{\text{uc}}=4\times 5$ and $N_{\text{uc}}=2\times 14$ in Fig. 4(e,f), respectively. A clear entanglement gap separating the low-lying PES levels from higher ones is identified. The number of PES levels below this gap exactly matches the typical counting of quasiparticle excitations resulting from the generalized Pauli principle of the non-Abelian Pfaffian or anti-Pfaffian state [25] (at most 2 particles in 4 consecutive orbitals). The particle-hole symmetry is explicitly broken by the non-uniform quantum geometries here, thus the particle-hole Pfaffian state [79] is less likely to be a competing phase compared to Pfaffian and anti-Pfaffian states [80]. The concrete nature of the non-Abelian state need more detailed examination by wave function overlap. From the correspondence between the PES and the quasiparticle excitations, the numerical results further suggest the non-Abelian nature of this half-filled state.
Discussions—The method studied here can be further generalized to Chern bands with band inversion resembling higher LL. Concretely, an ideal local band inversion is introduced with a simple $\mathbf{k}\cdot\mathbf{p}$ model [81]
$\mathcal{H}_{\text{local}}(\mathbf{k})=\begin{pmatrix}\alpha\mathbf{k}^{2}+\Delta&v_{F}k_{+}\\ v_{F}k_{-}&-\beta\mathbf{k}^{2}-\Delta\end{pmatrix}$, (7)
where $k_{\pm}=k_{x}\pm ik_{y}$. When $\Delta=-v_{F}^{2}/2(\alpha+\beta)$, both bands satisfy the trace and determinant conditions at any $\mathbf{k}$ with the quantum metric given by $g_{\mu\nu}(\mathbf{k})=(1/2)\left|\mathcal{B}(\mathbf{k})\right|\delta_{\mu\nu}$. Such an ideal local band inversion consistently yields a local Chern number $\mathcal{C}_{\text{local}}=1$ and a local average trace condition $\mathbb{T}_{\text{local}}=1$ when integrated over the region surrounding the inversion point. This framework explains why the orbitals $s$, $_{-}, p+p_{+}$ form a complete basis for describing the LLL and 1LL, as illustrated in Fig. 1 and Fig. 3. For the LCB, a single band inversion occurs between the $s$ and $p_{-}$ orbitals at the $\Gamma$ point. As a result, both $\mathcal{B}(\mathbf{k})$ and $\text{Tr}[g(\mathbf{k})]$ are concentrated around $\Gamma$, and the integral of them around $\Gamma$ yield a Chern number $\mathcal{C}=1$ and an average trace condition $\mathbb{T}\approx 1$. For the 1CB, two band inversions occur: one between $p_{-}$ and $s$ at $\Gamma$, and the other between $p_{-}$ and $p_{+}$ at $K$. The integral of $\mathcal{B}(\mathbf{k})$ around $\Gamma$ gives 1-1, and around $K$, it gives +1+1; the corresponding local $\mathbb{T}$ are both close to $1$. This results in a net Chern number $\mathcal{C}=2-1=1$, and a total trace condition $\mathbb{T}=\left|2\right|+\left|-1\right|=3$. This construction can be extended to higher Chern bands. As shown in Fig. 1(b,e), the band inversion at $M$ between $p_{+}$ and upper $d$ orbital contributes a local Chern number $\mathcal{C}_{\text{local}}=1$ and $\mathbb{T}_{\text{local}}=1$ at each $M$. Combined with the band inversion between $p_{+}$ and $p_{-}$ at $K$, this yields a total Chern number $\mathcal{C}=3-2=1$ and $\mathbb{T}=\left|3\right|+\left|-2\right|=5$ for the second Chern bands (i.e., the third band from the bottom), which matches the ideal value for the second LL. These results show that the quantum geometry of the constructed Chern bands closely matches that of generalized LL [73]. It is worth mentioning that the constructed Chern bands only have a perfect quantum weight at the penalty of Berry curvature flatness, since $\mathcal{B}(\mathbf{k})$ always concentrates at the local band inversion point. To obtain the FCI state in the partially filled Chern bands within exact diagonalization, a delicate balance between $\mathcal{B}(\mathbf{k})$ flatness and ideal $\mathbb{T}$ should be considered.
We also examine a minimal model that incorporates short-range hopping up to third-nearest neighbors, yielding a similar band structure and quantum geometry. Importantly, the many-body topological states remain robust in this setting [59], making it suitable to implementation in cold atom systems [82, 83, 84, 85], since only short-range density-density interactions are required. Optical lattices with synthetic gauge fields offer a controllable platform for realizing such orbital-based models with tunable interactions. Crucially, the lattice formulation is not a mere simplification but a central advantage, opening a practical route toward experimental realizations of fractional Chern insulators and non-Abelian phases in ultracold gases.
We note that the orbital construction of LL presented here is consistent with the concept of higher vortexability for the 1LL, as proposed in [40]. Specifically, the 1CB in our model is not self-vortexable, due to its large quantum metric trace $\mathbb{T}$. However, when the LCB and 1CB are considered together, we observe a nearly ideal combined trace $\mathbb{T}=2.2$ for the two lowest bands. This behavior originates from the band inversion between the $p_{-}$ and $p_{+}$ at the $K$ point. Our results thus offer a new perspective on understanding the vortexability of the lowest few LLs from a lattice orbital viewpoint.
Our orbital-based construction of LLs provides a complementary approach to QH lattice models. The well-known Kapit–Mueller (KM) model offers an exact lattice realization of LLL wavefunctions through engineered hopping terms [86, 87]. In contrast, our scheme relies on sequential band inversions among local Wannier orbitals of $s$, $_{-}, p+p_{+}$ character. This mechanism not only reproduces the LLL but also generates higher LLs within a unified multiband framework, thereby offering direct access to non-Abelian states inherent to higher levels.
Acknowledgements.
Acknowledgments—We thank Biao Lian and Dung-Hai Lee for helpful discussions. This work is supported by the Natural Science Foundation of China through Grants No. 12350404 and No. 12174066, the Innovation Program for Quantum Science and Technology through Grant No. 2021ZD0302600, the Science and Technology Commission of Shanghai Municipality under Grants No. 23JC1400600, No. 24LZ1400100 and No. 2019SHZDZX01, and is sponsored by the “Shuguang Program” supported by the Shanghai Education Development Foundation and Shanghai Municipal Education Commission. H.W. and R.S. contributed equally to this work.